# MetroGS: Efficient and Stable Reconstruction of Geometrically Accurate High-Fidelity Large-Scale Scenes

### v1

@@ 5.2 Baselines
We compare our method against a broad set of state-of-the-art surface reconstruction methods. For NeRF-based methods, we include NeuS [wang2021neus] and Neuralangelo [li2023neuralangelo]. For GS–based methods, we compare against SuGaR [guedon2024sugar], 2DGS [huang20242d], GOF [yu2024gaussian], CityGS [liu2024citygaussian], CityGS-$\mathcal{X}$ [gao2025citygs], and CityGSV2 [liu2024citygaussianv2]. We assess reconstruction quality from both visual and geometric perspectives. For large-scale scene reconstruction, we select CityGSV2 as a representative baseline, as it is among the best-performing open-source methods in terms of geometric reconstruction quality.
@@ 5.3 Main Results
As shown in Tab. 1, we compare our proposed method with several SOTA methods on the GauU-Scene [xiong2024gauu] dataset, which contains representative real-world urban scenes. The results show that our method achieves superior geometric reconstruction and rendering performance, ranking first on most metrics. Compared to CityGSV2 [liu2024citygaussianv2], our method improves PSNR by 0.88 dB on average and boosts the F1-score by 0.033, consistently outperforming it across all metrics. Tab. 2 presents the comparison results on the synthetic MatrixCity [li2023matrixcity] dataset. Our method again achieves the highest F1-score, with an average improvement of 0.11 over CityGSV2, indicating reliable and accurate geometric reconstruction across different scene types and data settings. In addition, even under the inherently stable illumination of synthetic data, our method maintains competitive rendering performance. Overall, these results validate the effectiveness of our method for robust, geometrically accurate, and high-fidelity large-scale scene reconstruction.
@@
To further validate the effectiveness of our method, we provide extensive visual comparisons. Fig. 4 shows image renderings and mesh reconstructions from synthetic MatrixCity [li2023matrixcity], where our method produces more accurate and more complete geometric reconstructions than CityGSV2 [liu2024citygaussianv2]. Fig. 5 presents qualitative comparisons of rendered images and corresponding depth maps from GauU-Scene [xiong2024gauu]. The first row depicts a scene under challenging lighting conditions: other methods suffer from floating artifacts caused by geometric errors, whereas our method yields more accurate illumination and more consistent geometric structures. The second and third rows contain scenes with rich texture details, where other methods show blurred or distorted structures in both the rendered images and depth maps. In contrast, our method preserves fine geometry and delivers visually coherent rendering results.
@@ 5.4 Ablation Studies
Table 3: Ablation on model components. The experiments are conducted on Russian Building scene of GauU-Scene [xiong2024gauu] dataset. We adopt a customized 2DGS [huang20242d] as our base method.
@@ 6 Conclusion
[…]
@@ Appendix A Implementation Details
For the GauU-Scene [xiong2024gauu] dataset, we conducted parallel training with a batch size of 4, targeting a total of 60,000 iterations. Subsequently, we train both the single-view and multi-view geometric refinement stages of $\mathcal{L}_{geo}$ for 30,000 iterations. During this process, $\lambda_{d}$ decreases from 0.5 to 0.005 as training progresses, while $\lambda_{n}$ is set to 0.0125, $\lambda_{s}$ to 0.1, and $\lambda_{mv}$ to 2.5. For $\mathcal{L}_{app}$, the weight $\lambda$ is set to 0.8. Densification terminates after the 15,000th iteration, with sparsity compensation parameters set to $S_{\text{th}}=20$ and $V_{\text{th}}=10$. The voxel size is set to 0.1 or 0.01 depending on the scale of the scene. For evaluation, only the view embeddings from the training set are available. Since the image filenames encode temporal information, we first use it to identify the two training views that are temporally closest to each test view. We select the candidate with the most similar camera pose to the test view. This nearest-neighbor assignment provides the interpolated view embedding for the test view.
@@
For the MatrixCity [li2023matrixcity] dataset, the Aerial and Street scenes were trained for 150,000 and 180,000 iterations, respectively. For $\mathcal{L}_{geo}$, single-view optimization is performed until the 50,000th iteration, followed by the switch to multi-view refinement. Densification is also terminated at the 50,000th iteration. All other training configurations follow those used for the GauU-Scene dataset. For evaluation, test image filenames lack temporal information, so each test view selects its most relevant training view solely based on camera-pose similarity. The corresponding view embedding is then used for image rendering.
@@
For geometric quality evaluation, we follow the parameter settings used in CityGSV2 [liu2024citygaussianv2]. Specifically, we render RGB images and depth maps from the training viewpoints and fuse them into a projected truncated signed distance functio (TSDF) volume [zeng20173dmatch] to extract surface meshes and point clouds. GauU-Scene uses a voxel size of 0.01, an SDF truncation of 0.04, and a depth truncation of 2.0. In MatrixCity, the Aerial split uses 0.01 / 0.04 / 5.0 for voxel size, SDF truncation, and depth truncation, respectively, whereas the Street split adopts 1 / 4 / 500.
@@
Table 4: Efficiency performance comparison on the GauU-Scene [xiong2024gauu] dataset. Entries marked with an asterisk (*) represent the intermediate results obtained after 30,000 training iterations.
@@
Table 5: Efficiency performance comparison on MatrixCity-Aerial [li2023matrixcity]. In CityGS-$\mathcal{X}$, which uses an anchor-based Gaussian representation, $\times 10$” denotes the Gaussians derived per anchor.
@@
Figure 8: Qualitative comparison of meshes on the GauU-Scene [xiong2024gauu] dataset. Our method achieves higher-quality results.
Figure 9: Mesh visualization comparison on MatrixCity-Aerial [xiong2024gauu]. Our method provides better results than the baselines.
Appendix B Hyperparameters of Other Methods
Appendix C Additional Results
C.1 Training Efficiency Analysis
Using a system with four RTX 3090 GPUs, we conducted a training efficiency comparison between CityGSV2 and CityGS-$\mathcal{X}$ on the GauU-Scene and MatrixCity-Aerial datasets, respectively. As shown in Tab. 4, our method consistently outperforms CityGSV2 in both rendering quality and geometric fidelity, while also demonstrating a significant improvement in training efficiency. Notably, even the intermediate results of our model at 30k iterations already surpass the final performance of CityGSV2, while requiring less than 25% of its training time. Across the GauU-Scene dataset, our final model achieves an average 2.55× training speedup relative to CityGSV2. Tab. 5 presents a comparison of training efficiency between CityGS-$\mathcal{X}$ and our method on the MatrixCity-Aerial dataset. Our approach achieves superior geometric fidelity (F1: 0.677 vs. 0.582) with a 1.7× reduction in training time, while maintaining comparable PSNR performance. Overall, these results highlight the remarkable speed and efficiency of our method. It is worth noting that CityGSV2 and CityGS-$\mathcal{X}$ adopt model-size reduction strategies such as trimming [fan2024trim] and anchor-based Gaussian compression [ren2024octree]. Enhancing model-size compactness therefore remains a promising direction for further improving the efficiency of our method.
Table 6: Quantitative results on the Mill19 [turki2022mega] dataset and UrbanScene3D [lin2022capturing] dataset. The best and second best results are highlighted. All missing results are denoted by a “–”.
@@ C.1 Training Efficiency Analysis
C.2 Additional Qualitative Comparison
Fig. 7 presents further visualization results for the ablation study. Our adopted pointmap assisted initialization effectively supplements sparse point cloud regions, thereby laying a solid geometric foundation for subsequent reconstruction. Progressive hybrid geometric refinement and depth-guided appearance modeling then collaboratively ensure the final geometric quality exhibits high accuracy and completeness.
In addition, we include more comprehensive qualitative comparisons with the baseline methods. Fig. 8 presents the mesh reconstruction visualization comparison on the GauU-Scene dataset. Given the relatively small size of the image data, we conducted an equivalent comparison in terms of training time: we trained our method for 30,000 iterations and compared its results with those of CityGSV2-coarse. The reconstructed meshes from our method are much cleaner, containing minimal spurious artifacts or floating mesh fragments. Fig 9 further presents a comparison of our method’s results against CityGSV2 and CityGS-$\mathcal{X}$ on the MatrixCity-Aerial dataset. The results indicate that our approach achieves a better balance between geometric accuracy and completeness.
C.3 Additional Dataset Evaluation
We have also conducted supplementary evaluations on the Mill19 [turki2022mega] and UrbanScene3D [lin2022capturing] datasets, which are widely used for assessing rendering quality in the field of large-scale scene reconstruction. Four scenes were selected: Building, Rubble, Residence, and Sci-Art. The configuration uses 100,000 training iterations, with 50,000 iterations allocated to each of the two geometric optimization stages. The densification process is terminated at the 30,000th iteration. The weight $\lambda_{s}$ set to 0.001. The remaining settings follow those used for GauU-Scene, as detailed in Sec. A.
@@ C.3 Additional Dataset Evaluation
Figure 10: Qualitative results on Mill-19 [turki2022mega] and Urbanscene3D [lin2022capturing] datasets. We compare against CityGS.
Appendix D Discussion
While our method successfully delivers efficient training, accurate geometry, and high rendering quality for large-scale scene reconstruction, it still presents the following limitations: Firstly, due to hardware constraints, memory consumption remains the primary bottleneck limiting the training scale, which to some extent weakens the model’s potential performance. Therefore, it is necessary to introduce techniques such as advanced pruning [mallick2024taming] and cache management [zhao2025clm] to mitigate memory challenges. Additionally, our method is based on 2DGS. Although it achieves excellent geometric reconstruction, its upper bound for rendering quality may still lag behind 3DGS. To address this, future work could consider introducing a new geometry representation similar to [jiang2025halogs] for complete decoupling of geometry and appearance to further realize improved geometric accuracy and rendering performance.

### v4

@@ 5.2 Baselines
We compare our method against a broad set of state-of-the-art surface reconstruction methods. For NeRF-based methods, we include NeuS [43] and Neuralangelo [27]. For GS–based methods, we compare against SuGaR [18], 2DGS [20], GOF [52], CityGS [33], CityGS-$\mathcal{X}$ [15], and CityGSV2 [34]. We assess reconstruction quality from both visual and geometric perspectives. For large-scale scene reconstruction, we select CityGSV2 as a representative baseline, as it is among the best-performing open-source methods in terms of geometric reconstruction quality.
@@ 5.3 Main Results
As shown in Tab. 1, we compare our proposed method with several SOTA methods on the GauU-Scene [48] dataset, which contains representative real-world urban scenes. The results show that our method achieves superior geometric reconstruction and rendering performance, ranking first on most metrics. Compared to CityGSV2 [34], our method improves PSNR by 0.88 dB on average and boosts the F1-score by 0.033, consistently outperforming it across all metrics. Tab. 2 presents the comparison results on the synthetic MatrixCity [24] dataset. Our method again achieves the highest F1-score, with an average improvement of 0.11 over CityGSV2, indicating reliable and accurate geometric reconstruction across different scene types and data settings. In addition, even under the inherently stable illumination of synthetic data, our method maintains competitive rendering performance. Overall, these results validate the effectiveness of our method for robust, geometrically accurate, and high-fidelity large-scale scene reconstruction.
@@
To further validate the effectiveness of our method, we provide extensive visual comparisons. Fig. 4 shows image renderings and mesh reconstructions from synthetic MatrixCity [24], where our method produces more accurate and more complete geometric reconstructions than CityGSV2 [34]. Fig. 5 presents qualitative comparisons of rendered images and corresponding depth maps from GauU-Scene [48]. The first row depicts a scene under challenging lighting conditions: other methods suffer from floating artifacts caused by geometric errors, whereas our method yields more accurate illumination and more consistent geometric structures. The second and third rows contain scenes with rich texture details, where other methods show blurred or distorted structures in both the rendered images and depth maps. In contrast, our method preserves fine geometry and delivers visually coherent rendering results.
@@ 5.4 Ablation Studies
Table 3: Ablation on model components. The experiments are conducted on Russian Building scene of GauU-Scene [48] dataset. We adopt a customized 2DGS [20] as our base method.
@@ 6 Conclusion
Acknowledgements
This work was in part supported by the Chinese Academy of Sciences, the Strategic Priority Research Program of the Chinese Academy of Sciences (XDA0450402), the Beijing Natural Science Foundation (L259015), the National Natural Science Foundation of China (62172392), and the Innovation Research Program of ICT CAS (E261070).
@@ Appendix A Implementation Details
For the GauU-Scene [48] dataset, we conducted parallel training with a batch size of 4, targeting a total of 60,000 iterations. Subsequently, we train both the single-view and multi-view geometric refinement stages of $\mathcal{L}_{geo}$ for 30,000 iterations. During this process, $\lambda_{d}$ decreases from 0.5 to 0.005 as training progresses, while $\lambda_{n}$ is set to 0.0125, $\lambda_{s}$ to 0.1, and $\lambda_{mv}$ to 2.5. For $\mathcal{L}_{app}$, the weight $\lambda$ is set to 0.8. Densification terminates after the 15,000th iteration, with sparsity compensation parameters set to $S_{\text{th}}=20$ and $V_{\text{th}}=10$. The voxel size is set to 0.1 or 0.01 depending on the scale of the scene. For evaluation, only the view embeddings from the training set are available. Since the image filenames encode temporal information, we first use it to identify the two training views that are temporally closest to each test view. We select the candidate with the most similar camera pose to the test view. This nearest-neighbor assignment provides the interpolated view embedding for the test view.
@@
For the MatrixCity [24] dataset, the Aerial and Street scenes were trained for 150,000 and 180,000 iterations, respectively. For $\mathcal{L}_{geo}$, single-view optimization is performed until the 50,000th iteration, followed by the switch to multi-view refinement. Densification is also terminated at the 50,000th iteration. All other training configurations follow those used for the GauU-Scene dataset. For evaluation, test image filenames lack temporal information, so each test view selects its most relevant training view solely based on camera-pose similarity. The corresponding view embedding is then used for image rendering.
@@
For geometric quality evaluation, we follow the parameter settings used in CityGSV2 [34]. Specifically, we render RGB images and depth maps from the training viewpoints and fuse them into a projected truncated signed distance functio (TSDF) volume [54] to extract surface meshes and point clouds. GauU-Scene uses a voxel size of 0.01, an SDF truncation of 0.04, and a depth truncation of 2.0. In MatrixCity, the Aerial split uses 0.01 / 0.04 / 5.0 for voxel size, SDF truncation, and depth truncation, respectively, whereas the Street split adopts 1 / 4 / 500.
Appendix B Hyperparameters of Other Methods
Appendix C Supplementary Method Description
Figure 7: Effectiveness of the appearance modeling.
C.1 Hybird Multi-View Refinement
During the multi-view consistency optimization stage, we first estimate an initial refined depth using Multi-View PatchMatch. The algorithm requires neighboring views. For each reference view, we preselect up to eight candidates based on camera orientation and spatial proximity, and randomly sample $K\geq 1$ views during multi-view optimization. Our implementation follows the multi-scale design of ACMM [49], but keeps the image resolution fixed while progressively reducing the patch size. Since PatchMatch operates on rendered depth, it is sensitive to local errors caused by floating Gaussians. To mitigate this issue, we enlarge the hypothesis sampling range and exclude the central region, adopting a non-local sampling strategy similar to DPE-MVS [7]. This design yields a reliable initial refined depth, even in large weakly textured regions.
Then we compute reprojection errors using neighboring rendered depths to filter out geometrically inconsistent estimates. Although effective, this step may remove correct depths due to occlusion or rendering noise. To recover such regions, we leverage monocular depth via local alignment. Specifically, we partition the image into non-overlapping image blocks and perform local alignment between the monocular depth and the filtered depth within each block (Eq. 8). Removed depths are restored if consistent with the aligned monocular depth. To improve robustness and completeness, we apply this alignment in a coarse-to-fine recursive manner with progressively smaller blocks. Finally, we obtain the desired depth map $\mathcal{D}_{mv}$, which is used to provide refined supervision for enforcing multi-view consistency.
C.2 Depth-Guided Appearance Modeling
Fig. 7 presents visual comparisons, demonstrating the effectiveness of appearance decoupling. To achieve this, we adopt a Tri-Mip structure to encode geometry-related features, which are queried in 3D space using depth information. The triplane representation ensures that the same spatial location yields consistent geometric features across different views. The mip mechanism further adapts features based on the distance to the current view, enabling consistent and accurate color mapping across varying scales and resolutions. In addition, we introduce view-specific appearance embeddings for each training view to avoid ambiguity in Gaussian geometry under appearance variations. This design incurs low computational overhead, is independent of the number of Gaussians, and can be removed after training.
@@ C.2 Depth-Guided Appearance Modeling
Table 4: Efficiency performance comparison on the GauU-Scene [48] dataset. Entries marked with an asterisk (*) represent the intermediate results obtained after 30,000 training iterations.
@@
Table 5: Efficiency performance comparison on MatrixCity-Aerial [24]. In CityGS-$\mathcal{X}$, which uses an anchor-based Gaussian representation, $\times 10$” denotes the Gaussians derived per anchor.
@@
Appendix D Additional Results
D.1 Training Efficiency Analysis
Using a system with four RTX 3090 GPUs, we conducted a training efficiency comparison between CityGSV2 and CityGS-$\mathcal{X}$ on the GauU-Scene and MatrixCity-Aerial datasets, respectively. As shown in Tab. 4, our method consistently outperforms CityGSV2 in both rendering quality and geometric fidelity, while also demonstrating a significant improvement in training efficiency. Notably, even the intermediate results of our model at 30k iterations already surpass the final performance of CityGSV2, while requiring less than 25% of its training time. Across the GauU-Scene dataset, our final model achieves an average 2.55× training speedup relative to CityGSV2. Tab. 5 presents a comparison of training efficiency between CityGS-$\mathcal{X}$ and our method on the MatrixCity-Aerial dataset. Our approach achieves superior geometric fidelity (F1: 0.677 vs. 0.582) with a 1.7× reduction in training time, while maintaining comparable PSNR performance. Overall, these results highlight the remarkable speed and efficiency of our method. It is worth noting that CityGSV2 and CityGS-$\mathcal{X}$ adopt model-size reduction strategies such as trimming [12] and anchor-based Gaussian compression [39]. Enhancing model-size compactness therefore remains a promising direction for further improving the efficiency of our method.
D.2 Impact of Prior Choices
As shown in Tab. 6 and the ablation studies in the main paper, incorporating priors leads to clear performance improvements, but the overall gains are not dominated by stronger priors. Further analysis reveals that the PointMap model contributes more significantly to the performance, highlighting the importance of high-quality initial points. In contrast, different monocular depth estimation methods have relatively limited impact on the final results.
Table 6: Ablation study on priors on the Russian Building [48]. Our method is robust to different choices of priors, with only limited impact on final performance.
Metrics Ours w/o priors VGGT + DAV2 VGGT + MoGe-2 PI3 + MoGe-2
PSNR / SSIM 24.85 / 0.807 24.94 / 0.812 24.93 / 0.812 24.94 / 0.814
F1-score 0.575 0.584 0.584 0.585
@@ D.2 Impact of Prior Choices
Figure 9: Qualitative comparison of meshes on the GauU-Scene [48] dataset. Our method achieves higher-quality results.
Figure 10: Mesh visualization comparison on MatrixCity [48]. Our method provides better results than the baselines.
D.3 Additional Qualitative Comparison
Fig. 8 presents visual results for further ablation study. Our adopted pointmap assisted initialization effectively supplements sparse point cloud regions, thereby laying a solid geometric foundation for subsequent reconstruction. Progressive hybrid geometric refinement and depth-guided appearance modeling then collaboratively ensure the final geometric quality exhibits high accuracy and completeness.
In addition, we include more comprehensive qualitative comparisons with the baseline methods. Fig. 9 presents the mesh reconstruction visualization comparison on the GauU-Scene dataset. Given the relatively small size of the image data, we conducted an equivalent comparison in terms of training time: we trained our method for 30,000 iterations and compared its results with those of CityGSV2-coarse. The reconstructed meshes from our method are much cleaner, containing minimal spurious artifacts or floating mesh fragments. Fig 10 further presents a comparison of our method’s results against CityGSV2 and CityGS-$\mathcal{X}$ on the MatrixCity dataset. The results indicate that our approach achieves a better balance between geometric accuracy and completeness.
Table 7: Quantitative results on the Mill19 [41] dataset and UrbanScene3D [31] dataset. The best and second best results are highlighted. All missing results are denoted by a “–”.
@@ D.3 Additional Qualitative Comparison
D.4 Additional Dataset Evaluation
We have also conducted supplementary evaluations on the Mill19 [41] and UrbanScene3D [31] datasets, which are widely used for assessing rendering quality in the field of large-scale scene reconstruction. Four scenes were selected: Building, Rubble, Residence, and Sci-Art. The configuration uses 100,000 training iterations, with 50,000 iterations allocated to each of the two geometric optimization stages. The densification process is terminated at the 30,000th iteration. The weight $\lambda_{s}$ set to 0.001. The remaining settings follow those used for GauU-Scene, as detailed in Sec. A.
@@ D.4 Additional Dataset Evaluation
Appendix E Limitation and Discussion
Although our method achieves efficient training, accurate geometry, and high-quality rendering for large-scale scene reconstruction, it still has several limitations. First, due to hardware constraints, memory consumption remains the primary bottleneck limiting the training scale, which in turn restricts the model’s performance upper bound. Therefore, it is necessary to incorporate techniques such as advanced pruning [35] and cache management [57] to alleviate memory pressure. In addition, our method is based on 2DGS. While it demonstrates strong performance in geometric reconstruction, its upper bound in rendering quality may still fall short of 3DGS. To address this, future work could explore introducing new geometric representations similar to [22] to achieve full decoupling of geometry and appearance, thereby further improving both geometric accuracy and rendering quality.
Figure 11: Qualitative results on Mill-19 [41] and Urbanscene3D [31] datasets. We compare against CityGS.
Furthermore, as shown in the mesh results of the street-view scene in Fig. 10, although our method achieves better overall performance, some limitations remain. Our results contain fewer floaters, indicating higher geometric accuracy, and benefit from pointmap to produce more complete reconstructions. However, training under street-view settings is more prone to exceeding GPU memory limits. A likely reason is that a large number of occluded Gaussians are still involved in rendering throughout the entire training process, leading to increased memory consumption. To ensure training stability, we introduce scale regularization, which, however, results in a limited number of large-scale Gaussians. This negatively affects the geometric quality of ground regions, leading to insufficient depth continuity and completeness. In addition, we adopt the median depth in 2DGS, which improves robustness but also compromises depth smoothness to some extent. These issues remain important directions for future improvement.